OpenStudy (maddiie):

Find the domain of each o the following rational functions. f(x)=3x-8 4x+20

OpenStudy (anonymous):

3x-8 domain = R (set of all real numbers 4x + 20 = R

OpenStudy (anonymous):

domain is the values x is allowed to have in the function. x can take on any value in either of those, so the domain is -infinity to +infinity, or all real numbers.

OpenStudy (maddiie):

So you aren't suppose to solve this?

OpenStudy (maddiie):

I like trying but i can't get this

OpenStudy (anonymous):

wait, is your question (3x-8)/(4x+20)? or 2 different questions?

OpenStudy (maddiie):

Yeah it's f(x)=3x-8 divided by 4x+20

OpenStudy (anonymous):

OK, so since you're not allowed to divide by zero, x can be anything EXCEPT the value that would make the denominator zero.

OpenStudy (anonymous):

so 4x + 20 = 0, solve for x, then the domain is "all x not equal to that number."

OpenStudy (maddiie):

Is it x=-5

OpenStudy (maddiie):

I'm not understanding

OpenStudy (anonymous):

right, so at x=-5 the function blows up. So the domain is anything BUT x = -5. See the graph: http://bit.ly/lBsh9u

OpenStudy (anonymous):

yes the number is -5

OpenStudy (maddiie):

That's it F(x)= -5

OpenStudy (anonymous):

the answer to "what is the domain" is "all x not equal to -5."

OpenStudy (maddiie):

So f(-5)= 3(-5)-8/-20+20 f=-3 huh

OpenStudy (anonymous):

I don't understand what you're asking / saying there. yes, f(-5) is undefined because of division by zero.

OpenStudy (maddiie):

I'm trying to figure it out and solve it

OpenStudy (anonymous):

figure what out? you already have the answer. The domain of a function is all values x is allowed to take in the function. In this case, x can take any value except those which cause division by zero. That's where x = -5. So the answer is "x can be anything except negative 5." You already figured it out!

OpenStudy (maddiie):

ohhhh lol oops